OpenStudy (anonymous):

What value of m makes the statement true? –(m + 15) = –18 + (–15) A. –33 B. –18 C. 3 D. 18

OpenStudy (anonymous):

@k-nowhere

OpenStudy (danjs):

you can solve the thing for M and see the value, or test each answer choice, and see if both sides remain =

OpenStudy (anonymous):

I agree^

OpenStudy (anonymous):

ok

OpenStudy (danjs):

–(m + 15) = –18 + (–15) add the right -(m + 15) = -33 multiply everything by -1 m + 15 = 33 move over the 15 m = 33 - 15

OpenStudy (anonymous):

so the answer is 33?

OpenStudy (danjs):

minus 15

OpenStudy (anonymous):

18

OpenStudy (anonymous):

Correct, 18.

OpenStudy (danjs):

yeah, put 18 in for m, and see if both sides are the same to check if you want

OpenStudy (anonymous):

thx